# TLDR: Unsupervised Goal-Conditioned RL via Temporal Distance-Aware Representations

### v1

@@ A.5 Evaluation Protocol
For Kitchen, we use the same $6$ single-task goal images used in LEXA [10], which consist of interactions with Kettle, Microwave, Light switch, Hinge cabinet, Slide cabinet, and Bottom burner. We report the total number of achieved tasks during evaluation.
@@
Appendix B More Ablation Studies
We conduct the ablation studies on the number of nearest neighbors $k$ (Figure 9) and $\dim\phi(\mathbf{s})$ (Figure 10) used in Equation 2. Figure 9 shows that different $k$ affects exploration in Ant, but for the other environments, the performance is not affected by the values of $k$. For $\dim\phi(\mathbf{s})$, the performance is nearly the same across different settings.
@@ Appendix B More Ablation Studies
(c) Humanoid-run
Figure 9: State coverage on state-based environments with different $k$. We measure the state coverage of our method with $k\in\{5,12,20\}$ used for calculating the TLDR reward in Equation 2. For Ant, $k=12$ works the best. For other environments, $k$ does not affect the state coverage.
@@
Appendix C More Qualitative Results
We include more qualitative results in Figures 14, 13, 11 and 12. For the qualitative results in Quadruped-Escape (Figure 12), we evenly select $48$ states satisfying $x^{2}+y^{2}=10^{2}$, where $x$, $y$ represents the agent position. The $z$ coordinate is selected as the minimum possible height that the agent does not collide with the terrain. For all environments, TLDR achieves the best goal-reaching behaviors compared to the other unsupervised GCRL methods, covering the goals in more diverse regions.
@@ Appendix C More Qualitative Results
Figure 12: Goal-reaching ability in Quadruped-Escape. We evaluate each method with the goals that are evenly selected at the same distance from the origin. TLDR can not only cover more regions but also have a better goal-reaching capability, compared to other methods.
@@
Figure 14: Goal-reaching ability in AntMaze-Ultra. TLDR can cover the most number of goals in AntMaze-Ultra compared to other methods.

### v2

@@ A.5 Evaluation Protocol
For Kitchen (Pixel), we use the same $6$ single-task goal images used in LEXA [10], which consist of interactions with Kettle, Microwave, Light switch, Hinge cabinet, Slide cabinet, and Bottom burner. We report the total number of achieved tasks during evaluation.
@@
For each environment, state coverage is calculated by the number of $1\times 1$-sized $(x,y)$-bins ($(x)$-bins for HalfCheetah) occupied by any of the training trajectories. For Kitchen (Pixel), the state coverage is calculated as the number of tasks achieved at least once during the last $100000$ environment steps.
Appendix B Sample Efficiency Comparison
We compare the sample efficiency of TLDR, METRA and PEG under the same setting as in Section 4.2. Since PEG requires longer training time for the same amount of environment steps compared to TLDR, PEG is trained for more than $72$ hours, while TLDR and METRA are each trained for $12$ hours.
Figures 9 and 10 illustrate the state coverage and goal-reaching metrics with respect to the number of environment steps used for training. TLDR exhibits superior state coverages and goal-reaching performance in hard-exploration tasks like AntMazes. In contrast, PEG tends to be more sample-efficient in environments with relatively low-dimensional state and action spaces, such as Ant and HalfCheetah, but it quickly converges to narrower regions in other environments that require hard-exploration (AntMazes) or have higher-dimensional state and action spaces (Quadruped-Escape, Humanoid-Run). METRA shows worse sample efficiency overall compared to TLDR.
PEG’s exploration via latent disagreement may over-prioritize less critical dimensions of the state spaces (e.g., joint angles) and may not scale well to high-dimensional observation spaces. Additionally, its reliance on expected temporal distances can be less effective for training a goal-reaching policy than TLDR’s optimal temporal distances, as shown in Section 4.4. Moreover, METRA’s skill learning objective can incentivize revisiting known distant states rather than exploring new ones, leading to suboptimal convergence.
Efficient exploration in high-dimensional spaces remains a major challenge in learning complex real-world tasks. Unlike other methods that quickly converge to suboptimal solutions in these settings, TLDR effectively handles this challenge and continues to improve steadily. We believe that increasing the update-to-data ratio or incorporating model-based reinforcement learning approaches could further enhance TLDR’s sample efficiency.
(a) Ant
(b) HalfCheetah
(c) Humanoid-Run
(d) Quadruped-Escape
(e) AntMaze-Large
(f) AntMaze-Ultra
[…]
(a) Ant
(b) HalfCheetah
(c) Humanoid-Run
(d) AntMaze-Large
(e) AntMaze-Ultra
Figure 10: Goal-reaching metrics of a goal-conditioned policy (sample efficiency). Similar to the results in Figure 9, while PEG can be trained efficiently in relatively low-dimensional tasks, TLDR has better sample complexity in more challenging tasks.
Appendix C More Ablation Studies
We conduct the ablation studies on the number of nearest neighbors $k$ (Figure 11) and $\dim\phi(\mathbf{s})$ (Figure 12) used in Equation 2. Figure 11 shows that in Ant environment, $k=12$ provides the best results, with exploration slightly degrading at $k=5$ or $20$; in the AntMaze-Large environment, the performance is rarely affected by the changes in $k$. Regarding $\dim\phi(\mathbf{s})$, the performance is nearly the same across different settings. Our main experimental results in Section 4.2 use $k=12$ and $\dim\phi(\mathbf{s})=4$, which demonstrates robust performance across diverse environments.
@@ Appendix C More Ablation Studies
Figure 11: State coverage on state-based environments with different $k$. We measure the state coverage of our method with $k\in\{5,12,20\}$ used for calculating the TLDR reward in Equation 2. For Ant, $k=12$ works the best. For AntMaze-Large, $k$ does not affect the final state coverage.
@@
Appendix D More Qualitative Results
We include more qualitative results in Figures 16, 15, 13 and 14. For the qualitative results in Quadruped-Escape (Figure 14), we evenly select $48$ states satisfying $x^{2}+y^{2}=10^{2}$, where $x$, $y$ represents the agent position. The $z$ coordinate is selected as the minimum possible height that the agent does not collide with the terrain. For all environments, TLDR achieves the best goal-reaching behaviors compared to the other unsupervised GCRL methods, covering the goals in more diverse regions.
@@ Appendix D More Qualitative Results
Figure 14: Goal-reaching ability in Quadruped-Escape. We evaluate each method with the goals that are evenly selected at the same distance from the origin. TLDR can not only cover more regions but also have a better goal-reaching capability than other methods.
@@
Figure 16: Goal-reaching ability in AntMaze-Ultra. Similar to Figure 15, TLDR can cover the most number of goals in AntMaze-Ultra, outperforming other methods.
Appendix E Unitree A1 Simulation Results
Figure 17: Unitree A1 Simulation Environment. To demonstrate the potential applicability for real-world robots, we choose the environment that simulates a Unitree A1 robot with 12 DoFs.
While most unsupervised goal-conditioned RL and skill discovery research currently focuses on simulated environments, unsupervised RL holds great potential for learning emergent and efficient skills for real-world robots.
To investigate whether TLDR can explore in environments with real-world robotic counterparts, we train TLDR on the Unitree A1 robot in simulation [44] (Figure 17), considering sim-to-real transfer approaches. TLDR and METRA are trained for $6$ hours, with the same hyperparameter settings we used in Quadruped-Escape and the episode length of $200$. For goal-conditioned evaluation, we sample goals with $(x,y)$-coordinates from $[-15,15]^{2}$.
As shown in Figure 18(a) and Figure 18(b), TLDR achieves substantially better state coverage and goal-reaching performance compared to METRA, suggesting the potential of TLDR for autonomous exploration and effective goal-reaching gaits learning in real robotics systems.
However, the learned behaviors with TLDR might be unsafe to transfer to reality since it does not impose any constraint on the learned behaviors beyond the goal-reaching objective. To address this, we test incorporating a safety reward for learning the exploration and goal-conditioned policies. The safety reward is defined as $r_{\text{safe}}=[0,0,1]\cdot\mathbf{v}_{\text{torso}}$, where $\mathbf{v}_{\text{torso}}$ is the orientation of robot torso, which equals to $[0,0,1]$ when the robot is upright.
The results in Figure 18 show that TLDR with this safety reward can match the performance of TLDR without regularization in terms of state coverage and goal-reaching metrics while also maximizing the safety reward. These findings indicate that TLDR is compatible with additional reward signals, and applying advanced safety-aware techinques [37, 38] could facilitate the learning of safer behaviors. Videos on learned behaviors can be found at https://heatz123.github.io/tldr.
(a) State Coverage
(b) Goal Distance
(c) Safety Reward
Figure 18: Learning curves of Unitree A1 Simulation Results. TLDR achieves better state coverage (a) and goal-reaching performance (b) compared to METRA. Since the learned behavior can be unsafe, we consider the setting that the safety reward (c) is given by $r_{\text{safe}}=[0,0,1]\cdot\mathbf{v_{torso}}$, where $\mathbf{v_{torso}}$ is the orientation of robot torso which equals to $[0,0,1]$ with upright direction. Even with adding the reward, TLDR can (a) still explore the state space and (b) learn effective goal-reaching behaviors (c) while maximizing the safety reward.
Appendix F Analysis on Pixel-based Environments
While achieving remarkable exploration and goal-reaching performance in state-based environments, exploration slows down in pixel-based environments, as observed in Figure 6. To identify the performance bottleneck for learning in pixel-based settings, we compare the performance when replacing pixel observations with state observations for the inputs of goal-conditioned policy, exploration policy, and TLDR encoder, respectively.
Figure 19 shows that the performance of TLDR becomes comparable to METRA when state observations are used instead of pixel observations for the goal-reaching policy, while other modifications do not improve upon the original TLDR. This suggests that the main bottleneck for exploration is likely to be the representations for the goal-conditioned policy. Based on this result, a promising future direction for improving TLDR in pixel-based environments could be the integration of advanced representation learning techniques [45, 46, 47] into our learning pipeline.
(a) State Coverage
(b) Goal Distance
Figure 19: Result of component-wise analysis in Quadruped (Pixel). To identify the bottleneck of exploration with pixel observations, we swap pixel observations to state observations for the input of goal-conditioned policy, exploration policy, and TLDR encoder, respectively. Exploration of TLDR significantly improves when we input state observations to the goal-conditioned policy, which suggests the main bottleneck for exploration is the representations for the goal-conditioned policy.
Additionally, while METRA quickly learns to achieve skills in Kitchen (Pixel) environment compared to TLDR, we observe that its performance degrades with continuous skills, as shown in Figure 20. This implies the difficulty of learning policies with continuous goals in TLDR’s goal-conditioned policy learning, possibly because learning to reach arbitrary states is more challenging than mastering a specific set of discrete behaviors. Investigating ways to restrict the size of the goal space in TLDR could be an interesting direction for future research.
Figure 20: Performance of METRA in Kitchen (Pixel) with different skill settings. When METRA uses continuous skill vectors in Kitchen (Pixel), METRA’s performance substantially degrades with continuous skills in Kitchen (Pixel) environment, which is similar to our setting of learning to reach any goals.
Appendix G Analysis on Goal-reaching Reward Design
We compare the goal-reaching performance of TLDR with different goal-conditioned policy learning methods using the same experimental setup as in Figure 8(a). As presented in Figure 21, other goal-reaching policy learning methods cannot reach the same level of performance as TLDR. This highlights the importance of our GCRL reward design on goal-reaching performance, consistent with the results in Figure 8(a), where TLDR achieves the best state coverage while other methods struggle.
(a) Ant
(b) AntMaze-Large
Figure 21: Goal-reaching metrics with GCRL reward design ablations. TLDR shows better goal-reaching performance compared to other choices of goal-conditioned policy learning methods, showing the effectiveness of our design of the GCRL reward.
To further isolate the impact of the exploration strategy and focus solely on goal-reaching policy learning, we evaluate the goal-reaching performance in an offline learning setting. In this setup, policies are trained on a fixed dataset of 1M samples collected from rollouts of a trained TLDR policy.
Although goal-reaching performances are degraded in this setting due to off-policy training, our choice of the goal-reaching reward still demonstrates superior results compared to other methods, as shown in Figure 22. This suggests that our design of the goal-reaching reward—minimizing the L2 distance to the goal in the temporal distance-aware representation space—provides effective signals for goal-reaching.
Figure 22: Goal-conditioned policy learning ablation with fixed dataset. With 1 million samples of rollouts from a trained TLDR policy, we learn the goal-conditioned policy without adding the data to the replay buffer, differing only in the goal-conditioned policy learning methods. TLDR shows the best performance in this setting where the impact of exploration is isolated.